OpenStudy (bekkah323):

find an equation of the line containing the given pair of points (1/4, -1/2) and (3/4, 6)

OpenStudy (bekkah323):

i got the slope as 1/2

OpenStudy (bekkah323):

nvm thats not the slope

OpenStudy (bahrom7893):

Well find the slope using: M = (y2-y1)/(x2-x1)

OpenStudy (luigi0210):

Do you know the formula?

OpenStudy (luigi0210):

My teacher bah has already provided it.. so nvm. Do you know how to USE it?

OpenStudy (luigi0210):

>.>

OpenStudy (bekkah323):

i think so but i feel like i got the wrong answer i got 13/2 over 1/2

OpenStudy (bekkah323):

so it would be 13

OpenStudy (luigi0210):

\[\frac{ \frac{ 4 }{ 6 }+\frac{ 1 }{ 2 } }{ 3-\frac{ 1 }{ 4 } }\]